Express as kg using decimals.
26 kg 50 g
Express as kg using decimals.
26 kg 50 g

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

We know that there are 1000 grams in 1 kg
26 kg 50 g = [(20 + 50 / 1000)] kg
= 26.050 kg

​Thanks
Shraddha Singh
3 years ago
Dear Student,
 
There are 1000 grams in 1 Kg.
So , we can say that,
 
1000 g = 1 Kg
1 g = 1/1000 Kg
50 g = 50/1000 Kg 
         = 0.050 Kg 
 
So, 26 Kg 50 g = (26 + 0.050) Kg
                           = 26.050 Kg